The following is a table of activities associated with a project at Rafay​ Ishfaq's software firm in​ Chicago, their durations and what activities each must​ precede:

Activity

Duration​ (weeks)

Precedes

A​ (start)

2

​B, C

B

2

E

C

6

F

E

3

F

F​ (end)

4

This exercise contains only parts​ b, c, and d.

​b) Define the critical​ path, listing all critical activities in chronological order.

Critical path activities are the activities that lie on the longest time path​ (in terms of​ time). The critical path follows the activities with slack​ = 0.

For A​ -C- E​ - F, the length is ____ weeks. ​(Enter your response as a whole number.​)

For A​ -B​- F, the length is ____ weeks. ​(Enter your response as a whole number.​)

The activities on the critical path are _____

​c) What is the project duration​ (in weeks)?

The total project completion time corresponds to the longest path in the​ network, which is ___ weeks. ​(Enter your response as a whole number.​)

Determine the slack time for each of the activities.

​(Enter your responses as whole numbers.​)

Activity

Slack Time​ (weeks)

A

B

C

E

F

What is the total slack for the non critical​ path(s) in the​ project?

There is only one noncritical path through the network.​ Thus, the slack​ (in weeks) associated with any and all noncritical paths through the project is

1111minus−1010equals=nothing

​week(s)

​(Enter your response as a whole number.​)
Answer #1

Network Diagram -

Possible Network Path and their Duration are -

ABEF --> 2+2+3+4 = 11 weeks
ACF --> 2+6+4 = 12 weeks

Critical path is the path with longest duration

Critical path = A-C-F

Project Duration = 12 weeks

Earliest Start (ES) = the earliest time when the activity can begin

Earliest Finish (EF) = ES + Activity Duration

Latest Finish (LF) = the latest time by whenthe activity can be finished without changing the project completion duration

Latest Start (LS) = LF - Activity Duration

Slack = LS-ES = LF-EF

Computing using the above gives us the below -

Activity Precedes Duration ES EF LS LF Slack
A B, C 2 0 2 0 2 0
B E 2 2 4 3 5 1
C F 6 2 8 2 8 0
E F 3 4 7 5 8 1
F - 4 8 12 8 12 0

Total slack for non critical activities = 1 + 1 = 2 weeks